# EEG2GAIT: A Hierarchical Graph Convolutional Network for EEG-based Gait Decoding

### v1

@@ 5.2 Ablation Analysis
The results of the loss ablation studies can be observed across the columns in Table 2. The HTSR loss, our proposed composite loss function combining frequency-domain loss, time-domain loss, and reward module, demonstrated superior performance compared to MSE loss across all baseline models and our EEG2GAIT model ($p\ll 0.01$ (Bonferroni corrected for 2 evaluated metrics and 11 models), paired-sample t-test, one-tail). Furthermore, as we progressively isolated each component of the loss function, the results showed that each individual component contributed positively to model performance for most models.
The HGP ablation study results can be found in the last two rows of tables in Table 1 and 2. When the HGP module was removed, the performance of EEG2GAIT declined. However, it still outperformed all baseline models, demonstrating the robustness and effectiveness of the core framework even without the HGP module.
@@ 5.3 Spatial Analysis
In addition to the performance metrics presented, we expanded our evaluation to include spatial feature importance analysis to better understand critical areas during decoding. To achieve this, we employed saliency mapping—a technique in machine learning that visualizes the importance of each input feature for the model’s predictions[simonyan2013deep]. This method highlights the input areas the model is most sensitive to when making predictions. The saliency map, $S$, is generated by calculating the gradient of the model’s output with respect to each input feature. The gradients are visualized to represent how variations in each input element, $X_{ij}$ (where $i,j$ are the spatial and temporal indices of $X$ , an input sample in $\mathbf{R}^{C\times T}$), influence the output prediction. The magnitude of each element $S_{ij}$ in $S$ illustrates the importance of the corresponding input pixel $X_{ij}$ to the output prediction.
@@
The saliency maps are plotted in Figure 4. The results indicated that the highlighted EEG channels during subjects’ walking were concentrated in the motor areas, especially in the lower-limb area, which is the anterior paracentral area[penfield1937somatic]. Our observations consistently pointed towards the motor areas, particularly channel Cz, FC1, FC2, Fz, CP1, CP2, exhibiting notably higher saliency compared to other regions. As shown in Figure 4, the average saliency maps across all subjects indicate strong role of central motor region in gait pattern encoding.
@@ 6.1 Hierarchical Graph Convolutional Network Pyramid
The proposed Hierarchical Graph Convolutional Network Pyramid (HGP) introduced a novel framework for capturing spatial relationships among EEG channels associated with gait-related brain activity by modeling EEG signals as a non-Euclidean structure. Direct comparison with existing state-of-the-art models in EEG-decoding revealed superior performance of the proposed approach, both in terms of achieved correlation and $R^{2}$-score between prediction and ground truth, and in the corresponding standard deviations (see Table 2). Importantly, the use of hierarchical graph convolutional layers with learnable adjacency matrices displayed the best regression performance of all compared methods, showing slight improvement compared to the basic EEG2Gait model without the hierarchical graph convolutional network (GCN) pyramid (HGP), specifically in the $R^{2}$-score. These findings stands in line with previous research, which documented effective motor activity decoding during gait by leveraging hierarchical graph convolutional layers with learnable adjacency matrices[sun2021adaptive, fu2021decoding].
@@ 6.3 Saliency Map Insights into Cortical Dynamics
Feature importance analysis using saliency maps revealed a consistent focus on sagittal-line electrodes in the frontocentral region, particularly around Cz, FC1, FC2, Fz, CP1, and CP2. These electrodes correspond to the bilateral leg motor area, aligning with prior findings on the neural representation of lower-limb motor control[zhang2017multiple, wei2023decoding]. The consistency of these results across participants underscores the neurophysiological relevance and generalizability of the extracted features, confirming the model’s robustness against individual variations in head physiology.

### v2

@@ 5.2 Ablation Analysis
The results of the loss ablation studies can be observed across the columns in Table 1 and 2. To assess statistical significance, we tested whether the performance improvements of HTSR loss over MSE loss were greater than zero. Results confirmed that HTSR loss produced significantly higher performance across all baseline models and our EEG2GAIT model ($p\ll 0.001$ for all three metrics, one-tailed paired-sample t-test). Furthermore, as we progressively isolated each component of the loss function, the results showed that each individual component contributed positively to model performance for most models.
To evaluate the contribution of the HGP module, we conducted an ablation study comparing EEG2GAIT with and without HGP. On the GPP dataset, the inclusion of HGP significantly improved performance across all metrics ($p\ll 0.001$). On the MoBI dataset, the improvements were also statistically significant, with $p=3.34\times 10^{-3}$ for $r$, and $p=1.13\times 10^{-2}$ for $R^{2}$ and $p=3.42\times 10^{-2}$ for MAE. Even without HGP, EEG2GAIT still outperformed all baseline models, highlighting the robustness of the core framework, while the inclusion of HGP provided consistent and statistically significant gains.
5.3 Impact of ICA-based Artifact Removal
We trained all models on the ICA-cleaned datasets following the same procedures described in Section 4.1. The results are summarized in Table 4. EEG2GAIT achieved a high decoding accuracy after ICA cleaning, with an average $r$ value of 0.935, slightly lower than its performance on the unprocessed data.
For baseline methods, the effect of ICA cleaning was mixed: several models (e.g., ContraWR, FFCL, TSception, SPaRCNet) showed performance gains after ICA, whereas others (e.g., ST-Transformer, EEGConformer, EEGNet, deepConvNet) exhibited decreases.
Table 4: Performance comparison on GED and MoBI datasets with and without ICA-based artifact removal. Values are reported as mean across participants. For each metric, the single best value across all models and ICA settings is highlighted in bold. Within each model, the better of w/o ICA vs w/ ICA is marked with . Values are presented in the order: without ICA / with ICA.
Model GED MoBI
$r$ value $R^{2}$ score MAE ($\downarrow$) $r$ value $R^{2}$ score MAE ($\downarrow$)
ContraWR [52] 0.6076 / 0.6263 0.3058 / 0.3142 0.5875 / 0.5698 0.4438 / 0.3140 -0.0021 / 0.0427 7.74 / 7.60
FFCL [27] 0.6473 / 0.6710 0.3666 / 0.4046 0.5579 / 0.5563 0.4460 / 0.4407 0.0337 / 0.1413 7.65 / 7.23
TSception [11] 0.7214 / 0.7257 0.4560 / 0.4813 0.5592 / 0.5451 0.5678 / 0.5527 0.2342 / 0.1918 7.11 / 7.33
TCN [18] 0.7335 / 0.7034 0.5244 / 0.4914 0.4887 / 0.5351 0.6040 / 0.6099 0.3492 / 0.3423 6.19 / 6.37
ST-Transformer [42] 0.9193 / 0.8898 0.8219 / 0.7681 0.2630 / 0.3153 0.7171 / 0.6760 0.4851 / 0.4099 4.94 / 5.92
EEGConformer [43] 0.9001 / 0.8739 0.7901 / 0.7597 0.2925 / 0.3226 0.6844 / 0.6027 0.4571 / 0.3325 5.35 / 6.27
SPaRCNet [20] 0.8852 / 0.8877 0.7532 / 0.7534 0.3077 / 0.3120 0.6838 / 0.6459 0.4188 / 0.3586 5.37 / 6.38
EEGNet [25] 0.8885 / 0.8641 0.7604 / 0.7150 0.3515 / 0.3642 0.6898 / 0.6495 0.4574 / 0.3988 5.71 / 5.74
deepConvNet [38] 0.9497 / 0.9224 0.8769 / 0.8375 0.2125 / 0.2504 0.7588 / 0.6854 0.5585 / 0.4539 4.63 / 5.49
EEG2GAIT (w/o HGP) 0.9526 / 0.9257 0.9008 / 0.8487 0.2056 / 0.2183 0.7629 / 0.7009 0.5720 / 0.4517 4.51 / 5.61
EEG2GAIT 0.9588 / 0.9351 0.9141 / 0.8651 0.1930 / 0.2107 0.7793 / 0.7126 0.5965 / 0.4835 4.38 / 4.85
@@ 5.4 Spatial Analysis
In addition to the performance metrics presented, we expanded our evaluation to include spatial feature importance analysis to better understand critical areas during decoding. To achieve this, we employed saliency mapping—a technique in machine learning that visualizes the importance of each input feature for the model’s predictions [41]. This method highlights the input areas the model is most sensitive to when making predictions. The saliency map, $S$, is generated by calculating the gradient of the model’s output with respect to each input feature. The gradients are visualized to represent how variations in each input element, $X_{ij}$ (where $i,j$ are the spatial and temporal indices of $X$ , an input sample in $\mathbf{R}^{C\times T}$), influence the output prediction. The magnitude of each element $S_{ij}$ in $S$ illustrates the importance of the corresponding input pixel $X_{ij}$ to the output prediction.
@@
The saliency maps are plotted in Figure 4. The results indicated that the highlighted EEG channels during subjects’ walking were concentrated in the motor areas, especially in the lower-limb area, which is the anterior paracentral area [31]. Our observations consistently pointed towards the motor areas, particularly channel Cz, FC1, FC2, Fz, CP1, CP2, CPz, exhibiting notably higher saliency compared to other regions. As shown in Figure 4, the average saliency maps across all subjects indicate strong role of central motor region in gait pattern encoding.
@@ 6.1 Hierarchical Graph Convolutional Network Pyramid
The proposed Hierarchical Graph Convolutional Network Pyramid (HGP) introduced a novel framework for capturing spatial relationships among EEG channels associated with gait-related brain activity by modeling EEG signals as a non-Euclidean structure. Direct comparison with existing state-of-the-art models in EEG-decoding revealed superior performance of the proposed approach, both in terms of achieved correlation and $R^{2}$-score between prediction and ground truth, and in the corresponding standard deviations (see Table 2). Importantly, the use of hierarchical graph convolutional layers with learnable adjacency matrices displayed the best regression performance of all compared methods, showing slight improvement compared to the basic EEG2Gait model without the hierarchical graph convolutional network (GCN) pyramid (HGP), specifically in the $R^{2}$-score. These findings stands in line with previous research, which documented effective motor activity decoding during gait by leveraging hierarchical graph convolutional layers with learnable adjacency matrices [45, 13].
@@ 6.2 Combined temporal and spectral losses
6.3 On the Role of ICA in Gait Decoding
Our control experiment with ICA-based artifact removal revealed a trade-off between artifact suppression and neural information preservation. While ICA effectively reduced motion- and eye-related artifacts, performance decreases in stronger models such as EEG2GAIT suggest that task-relevant neural components were inadvertently removed. On average, 8–10 components (over 15% of total) were discarded, which partly explains the diminished effectiveness of ICA for high-performing models.
Since ICA requires offline decomposition and manual or semi-automatic component selection, it is not directly compatible with real-time BCI systems. Our results indicate that EEG2GAIT achieves high decoding accuracy under online-compatible preprocessing (i.e., band-pass filtering) without relying on ICA. This suggests that retaining neural information—even at the cost of some residual artifacts—benefits gait decoding more than aggressive artifact removal. Consequently, we advocate band-pass filtering alone for real-time applications, balancing artifact suppression and information retention.
@@ 6.4 Saliency Map Insights into Cortical Dynamics
Feature importance analysis using saliency maps revealed a consistent focus on sagittal-line electrodes in the frontocentral region, particularly around Cz, FC1, FC2, Fz, CPz, CP1, and CP2. These electrodes correspond to the bilateral leg motor area, aligning with prior findings on the neural representation of lower-limb motor control [56, 51]. The consistency of these results across participants underscores the neurophysiological relevance and generalizability of the extracted features, confirming the model’s robustness against individual variations in head physiology.
6.5 Practical Applications in Assistive Technologies
Beyond benchmarking performance, the proposed EEG2GAIT framework holds considerable potential for practical applications in assistive technologies. First, our preliminary experiments suggest that variations between EEG-predicted gait dynamics and actual recorded trajectories are associated with the rehabilitation stages of stroke patients. This suggests that EEG2GAIT could serve as a fine-grained tool for evaluating motor recovery, thereby enabling objective assessment of rehabilitation progress and providing clinicians with a quantitative biomarker for monitoring patient-specific recovery trajectories.
Second, EEG2GAIT’s capability to predict lower-limb joint angles with high temporal resolution enables precise identification of gait phases, particularly the swing phase. Accurate detection of these critical gait events is especially important in rehabilitation, as it allows clinicians and assistive devices to pinpoint problematic intervals within the gait cycle. Importantly, this temporal precision permits targeted neural modulation during specific phases of dynamic movement, supporting phase-specific interventions that can enhance motor recovery. Additionally, it facilitates real-time control of lower-limb exoskeletons and prosthetic devices, where timely recognition of motor intent improves synchronization between the user and the device, thereby enhancing both safety and functional effectiveness in brain-controlled assistive systems.
6.6 Limitations
Despite its promising performance, EEG2GAIT has two main limitations. First, most practical applications of brain-controlled exoskeletons are based on motor imagery, whereas our current model relies on decoding EEG signals recorded during actual walking. This gap may limit direct applicability in scenarios where overt movement is not possible.
Second, EEG2GAIT is currently a subject-dependent model, which means that its generalization to unseen individuals is limited. Future work will explore subject-independent or transfer learning strategies to improve cross-subject generalizability and reduce calibration requirements.
6.7 Future Work
To address the identified limitations and advance the clinical applicability of EEG2GAIT, several research directions will be pursued in future work.
Cross-Subject Generalization: Our analysis of saliency maps has revealed important insights that will guide the development of subject-independent models. While we observed consistent patterns across most subjects, where brain regions associated with lower-limb movement consistently showed higher saliency than other areas, each subject still exhibits distinct saliency map characteristics. Notably, we discovered that subjects’ saliency maps can be clustered, indicating that some subjects share higher similarity in their neural activation patterns than others.
Based on these findings, saliency map-based transfer learning between subjects represents a promising optimization direction for developing generalized models. We plan to develop a hierarchical transfer learning framework that: (1) identifies subject clusters based on saliency map similarity using unsupervised learning algorithms, (2) applies cluster-specific fine-tuning strategies to new subjects, and (3) incorporates few-shot learning techniques to minimize calibration data requirements. This approach could significantly reduce the calibration time required for new users while maintaining prediction accuracy.
Motor Imagery Integration: Future work will explore bridging the gap between actual walking and motor imagery by investigating shared neural representations between these two modalities. We will develop domain adaptation techniques that can transfer knowledge learned from actual walking data to motor imagery scenarios, enabling the system to work with patients who cannot perform overt movements.
Clinical Translation: From a clinical perspective, we will focus on validating the model with patients having different neurological conditions (stroke, spinal cord injury, etc.) and developing adaptive algorithms that can adjust to disease progression or recovery. Additionally, we will integrate EEG2GAIT with existing rehabilitation protocols and assistive devices to demonstrate its practical utility in real-world therapeutic settings.